Metlock, Inc. has 10,200 shares of 9 %, $ 100 par value, cumulative preferred stock outstanding at December 31, 2022. No dividends were declared in 2020 or 2021

If Metlock wants to pay $ 375,000 of dividends in 2022 , what amount of dividends will common stockholders receive?

Dividends allocated to common stock $ _______ 
Answer #1

Preferred dividends per year = 10,200*9%*100 = 91,800

Dividends allocated to common stock

= Net income - 3 years of preferred stock dividends

= 375,000 - (91,800*3)

= 99,600

Comment if you face any issues instead of downvoting